OpenStudy (anonymous):

Which is the radius of a circle whose equation is (x + 1)2 + (y - 1)2 = 25? 30 25 5 4

OpenStudy (zarkon):

\[(x-a)^2+(y-b)^2=r^2\] equation of a circle centered at (a,b) with radius r

OpenStudy (anonymous):

So 25=r^2 what is r=

OpenStudy (anonymous):

Huh? can you write the steps out so i know how to Do it?

myininaya (myininaya):

\[(x-(-1))^2+(y-1)^2=5^2\] using what zarkon wrote what is a=? what is b? what is r?